I need help to solve this problems.
-6=2/3z
5/9p=-10

Answer by checkley77(12844)

-6=2/3z
2Z/3=-6
2Z=-6*3
2Z=-18
Z=-18/2
Z=-9 ANS.
-------------
5/9p=-10
5P/9=10
5P=10*9
5P=90
P=90/5
P=18 ANS.